OpenStudy (anonymous):

help me @Daish96 @dumbsearch2 @damien @Danny_Boy

OpenStudy (anonymous):

@countonme123

OpenStudy (anonymous):

omgggggggggg idk

OpenStudy (anonymous):

not in that problem

OpenStudy (anonymous):

@damien @Daish96 @DLS @dietrich_harmon @Koikkara @kugler97 @Kpinky

OpenStudy (koikkara):

@InYourHead .....

OpenStudy (anonymous):

@ThatHippieKid @haleylou56 @heena @hewsmike @hannahbabiexoxo

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

@nincompoop @NickR @nolesgirl328 @nolesgirl328

OpenStudy (anonymous):

64b^10/b^9c^5= 64b/c^5

OpenStudy (anonymous):

omggggggg confused thats the anwser?????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ummm thanks can u help me with these other 3 questions?

OpenStudy (anonymous):

but on one condition u hav to also try with me and have to try to undrstand how towork on such question is it fine?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

)Robert loves football and wants to start practice throwing daily. The first day he practices for 3 minutes. The next day he triples his time. Robert continues to triple his practice time every day. How many minutes did Robert practice on the fourth day?

OpenStudy (anonymous):

ok on firt day he did 3^1-3min den on second day he triples his time mean 3^3=3x3x3=27min

OpenStudy (anonymous):

now on third day u tell me

OpenStudy (anonymous):

27x3=81

OpenStudy (anonymous):

i think triples mean increasing power means 27^3

OpenStudy (anonymous):

danya what happen?

OpenStudy (anonymous):

27x27x27=19683

OpenStudy (anonymous):

???????????????????????

OpenStudy (anonymous):

@heena ??????????????????

OpenStudy (anonymous):

what dear ? now work for fourth dat imilarly as u did for third day

OpenStudy (anonymous):

similaraly

OpenStudy (anonymous):

didnt u get the trick how to work on it?

OpenStudy (anonymous):

7625597484987

OpenStudy (anonymous):

is that right???

OpenStudy (anonymous):

@heena

OpenStudy (anonymous):

@jonnymiller @Jenn777 @hewsmike @heena @whpalmer4 @Agent_Sniffles @ash2326 @Koikkara @kelliegirl33 is this right

OpenStudy (anonymous):

@ash2326 i think i made a mistake help her plz

OpenStudy (anonymous):

orry danya but i think i m not correct here

OpenStudy (anonymous):

Im so lost lol

OpenStudy (anonymous):

the questiopn is at the top

OpenStudy (anonymous):

its okay @heena

OpenStudy (ash2326):

@danya1 Always ask new question as a new post Robert loves football and wants to start practice throwing daily. The first day he practices for 3 minutes. The next day he triples his time. Robert continues to triple his practice time every day. How many minutes did Robert practice on the fourth day?

OpenStudy (ash2326):

He triples his time means times 3 First day 3 minutes=3 mins Second 3*3 minutes=9 mins Can you find for the third day and 4th day? It's just multiplication :)

OpenStudy (anonymous):

yes thats the question

OpenStudy (anonymous):

do i do 9x3??

OpenStudy (anonymous):

Is this how to do it.....are you sure it doesn't have anything to do with exonents ? first day = 3 min 2nd day = 3 x 3 = 9 3rd day = 9 x 3 = 27 4th day = 27 x 3 = 81

OpenStudy (anonymous):

oops...typo...exponents ?

OpenStudy (ash2326):

yes @danya1

OpenStudy (anonymous):

or is it 9x9x9

OpenStudy (ash2326):

9*3

OpenStudy (anonymous):

see...that is what is confusing me....exponents

OpenStudy (anonymous):

I suppose ash is right.....we are making this too hard when it isn't.

OpenStudy (ash2326):

thrice is 3 times

OpenStudy (anonymous):

Thanks for explaining ash....oh, and I am sorry I stepped on your turf...I will do better

OpenStudy (ash2326):

no problem @kelliegirl33

OpenStudy (anonymous):

you deserve the medal ash :)

OpenStudy (ash2326):

Do you get @danya1

OpenStudy (anonymous):

so the second day was 9 and the third is 9x3=27 and the fourth day is 27x3=81

OpenStudy (ash2326):

Yes @danya1 :D

OpenStudy (anonymous):

ash is good

OpenStudy (ash2326):

Thanks @kelliegirl33 :) Likewise :D

OpenStudy (anonymous):

can u help me withe the first problem i wrote at the top thanks

OpenStudy (ash2326):

Okay, you want to simplify it \[(\frac{2b^2}{b^4c})^5\] First cancel the common terms \[(\frac{2\cancel{b^2}}{b^{\cancel 4 2}c})^5\] so we get \[(\frac{2}{b^2c})^5\] Do you understand till here?

OpenStudy (anonymous):

ummmmmmmmmmmmmm yea

OpenStudy (ash2326):

Now just raise the terms to the power \[\frac{2^5}{b^{2^5}c^5}\] \[\frac{32}{b^{10}c^5}\]

OpenStudy (anonymous):

ok thats it

OpenStudy (ash2326):

yes

OpenStudy (anonymous):

thankswhat about this one

OpenStudy (ash2326):

@danya1 Please close this question and ask this question in a new post :) That way you'll get help from many users :)